OpenStudy (anonymous):

tan^2 x / sec x+1 I need to get it down to one function with no fractions

OpenStudy (anonymous):

\[\frac{\tan^2x}{\sec x+1}\cdot\frac{\sec x-1}{\sec x-1}\] \[\frac{(\tan^2x)(\sec x-1)}{\sec^2 x-1}\] Use an identity to rewrite the denominator.

OpenStudy (anonymous):

where do I go from there then? I don't see how things can cancel out or go away

OpenStudy (anonymous):

\[\sec^2x=\tan^2x+1,\text{ right?}\\ \frac{(\tan^2x)(\sec x-1)}{\tan^2x}\\ \sec x-1\]

OpenStudy (anonymous):

I shouldn't have said anything about rewriting the denominator, sorry.

OpenStudy (anonymous):

no you have to rewrite the denominator to get it to be a minus 1 instead of plus, thank you so much!